simplify
(6^(x-5)) / (6^(3-x))

Answer by fractalier(2101)

Okay, we we divide like bases, we subtract the exponents, so we get
(6^(x-5)) / (6^(3-x)) =
6^ ((x-5) - (3-x)) =
6^(2x-8) =
6^(2(x-4)) =
36^(x-4)